# R-SFLLM: Jamming Resilient Framework for Split Federated Learning with Large Language Models

### v2

@@ Abstract
Split federated learning (SFL) is a compute-efficient paradigm in distributed machine learning (ML), where components of large ML models are outsourced to remote servers. A significant challenge in SFL, particularly when deployed over wireless channels, is the susceptibility of transmitted model parameters to adversarial jamming that could jeopardize the learning process. This is particularly pronounced for word embedding parameters in large language models (LLMs), which are crucial for language understanding. In this paper, rigorous insights are provided into the influence of jamming LLM word embeddings in SFL by deriving an expression for the ML training loss divergence and showing that it is upper-bounded by the mean squared error (MSE). Based on this analysis, a physical layer framework is developed for resilient SFL with LLMs (R-SFLLM) over wireless networks. R-SFLLM leverages wireless sensing data to gather information on the jamming directions-of-arrival (DoAs) for the purpose of devising a novel, sensing-assisted anti-jamming strategy while jointly optimizing beamforming, user scheduling, and resource allocation. Extensive experiments using BERT and RoBERTa models demonstrate R-SFLLM’s effectiveness, achieving close-to-baseline performance across various natural language processing (NLP) tasks and datasets. The proposed methodology further introduces an adversarial training component, where controlled noise exposure significantly enhances the LLM’s resilience to perturbed parameters during training. The results show that more noise-sensitive models, such as RoBERTa, benefit from this feature, especially when resource allocation is unfair. It is also shown that worst-case jamming in particular translates into worst-case model outcomes, thereby necessitating the need for jamming-resilient SFL protocols.
Index Terms:
6G, anti-jamming, large language model, resource allocation, split federated learning, wireless sensing
@@ I Introduction and Motivation
Future 6G networks are anticipated to introduce a substantial leap toward highly integrated and intelligent connectivity at the edge, enabled by artificial intelligence (AI) [1] and machine learning (ML) [2]. However, in this envisioned hyper-connected and AI-assisted network, important questions and stringent requirements on resilience and trustworthiness arise, both from a user and network perspective [3]. This imposes significant design challenges for emerging technologies, such as distributed and collaborative ML (DCML) [4], which may be targeted by adversarial attacks over the wireless medium. For instance, the distributed training of large language models (LLMs) faces unique challenges in ensuring data integrity and model robustness due to highly sensitive word embedding parameters. Much research has focused on addressing these challenges via adversarial AI training methods and model-based solutions [5, 6], and in [7], the authors closely investigated such attack and defense strategies in 6G networks, exposing critical vulnerabilities across all network layers. However, a number of important research questions remain underexplored, such as how these attacks can be orchestrated in practice, and how current and next-generation wireless network architectures, systems, and technologies can introduce proactive defense mechanisms, ideally by-design.
I-A Adversarial Poisoning in Wireless Federated LLM Training
@@
Motivated by the increasing importance of end-user data privacy and associated privacy protection laws [8], DCML has been gradually shifting toward mechanisms such as federated learning (FL) [9] and split FL (SFL) [10]. These have proven to be privacy-preserving as only the respective model parameters, some parts of it or model gradients need to be exchanged. SFL in particular has emerged as a compute-efficient federated protocol, suitable for distributed training of large ML model architectures. Unlike traditional FL, in which the entire model is trained on each client, SFL splits the model, allowing for more compute-intensive parts to be outsourced to a remote server. This approach is particularly advantageous for LLM architectures that cannot be entirely processed at resource-constrained edge devices due to computational and memory limitations [11]. However, the practical realization of (S)FL over wireless networks faces challenges from the inherently unreliable wireless medium, limited bandwidth, and suboptimal resource allocation [12]. In addition, malicious actors such as jammers could target privacy and security aspects of (S)FL systems by intentionally poisoning data and models through adversarial noise. The particular importance of studying adversarial jamming attacks in SFL with LLMs is motivated by recent results from natural language processing (NLP) research [13, 14]. The authors in [13] study the susceptibility of LLMs to word embedding poisoning caused by noisy perturbations and show that by altering even a single word embedding vector, an adversary can subtly manipulate a model to react abnormally to specific trigger words. Moreover, the severity of word embedding poisoning attacks for federated networks was studied in [14], showing that even a small number of compromised clients is sufficient to effectively deteriorate the global model. In federated systems over wireless networks, adversarial jamming emerges as a realistic threat for orchestrating such poisoning attacks by corrupting sensitive word embeddings during transmission. This covert adversarial perspective is in contrast to prior works on jamming in FL [15, 16], which do not consider the implications of poisoning the model’s reasoning capabilities. This is particularly critical in SFL, where word embeddings might be directly transmitted as intermediate split parameters. However, the detrimental impact of such attacks in SFL remains underexplored.
I-B Proactive and Resilient-by-Design Anti-Jamming in SFL
@@
To preemptively safeguard LLM parameter transmissions in SFL against adversarial jamming attacks, proactive and in particular resilient-by-design approaches are needed [17]. This requires a simultaneous co-design of AI, resilience, beamforming, user scheduling, and resource allocation, thereby integrating resilience from a bottom-up approach, by design. In addition, such proactive defense mechanisms need to be agnostic of the respective jamming capabilities, including physical and spatial features. This is not the case for some more recent works, which tend to impose strong assumptions on the adversary’s knowledge and setup. For instance, [18] and [19] only consider single- or few-antenna jammers with common secrets being exchanged between legitimate parties. This essentially excludes so-called worst-case jammers with extensive system knowledge and capabilities [20, 21]. In order to develop universally applicable defense strategies for a wide range of jamming scenarios in SFL, system performance needs to be guaranteed for the worst-case. Thus, we need to generalize toward intelligent and reconfigurable worst-case jammers. In our prior work in [22], we motivated the study of how sensing-assisted network information can be harnessed to enhance existing mitigation schemes without the need for otherwise precise jamming statistics. Therein, we have shown that information on the jamming signal directions-of-arrival (DoAs) can be used to devise MIMO-OFDM anti-jamming strategies with exceptional performance. However, we did not discuss whether such sensing-assisted defense strategies can be straightforwardly applied to enhance the resilience in SFL over wireless networks. In particular, the impact of worst-case jamming needs to be quantified in order to study its influence on LLM model poisoning as compared to conventional jammers. In the subsequent sections, we provide thorough insights on these aspects, including an analysis on the minimum system rate that guarantees a reliable and resilient SFL training.
I-C Contributions
@@
The main contribution of this paper is an analysis and framework for resilient SFL over wireless networks that will help close the gap between adversarial jamming attacks in SFL and LLM model poisoning. We provide insights into how jamming LLM word embeddings affects the global model training and how the latter can be efficiently safeguarded by MIMO signal processing at the availability of sensing-assisted information. In summary, our key contributions include:
@@
  • We derive an analytical expression for the ML training loss divergence based on a relaxed $(\boldsymbol{L}_{0},\boldsymbol{L}_{1})$-smoothness assumption for LLM transformer architectures in the case of corrupted word embeddings. We show that its upper bound depends on the communication mean squared error (MSE), thereby motivating a wireless approach to resilience in SFL.
  • We provide a novel analysis on the minimum system rate which guarantees a robust and reliable SFL training over the wireless network, thereby characterizing minimum network conditions based on the outage rate caused by the jammer.
@@
  • We develop R-SFLLM, a novel, sensing-assisted anti-jamming framework for resilient SFL with LLMs, which leverages the jamming signal’s DoAs to devise an anti-jamming strategy formulated as a joint optimization problem for beamforming, user scheduling, and resource allocation while maximizing the sum rate of the SFL participants. In this problem, any explicit knowledge about the jamming statistics is replaced by a surrogate expression that depends only on the jamming DoAs. We provide an efficient solution to the problem using an iterative water-filling approach [23].
  • In order to benchmark R-SFLLM against worst-case conditions in SFL, we utilize the worst-case jamming strategy in our prior work [22], which minimizes the sum rate instead.
@@
  • We provide extensive simulations for BERT and RoBERTa models on various datasets, demonstrating near-optimal performance when anti-jamming is enabled and significantly worse outcomes for unprotected scenarios. We show that R-SFLLM introduces an additional adversarial training component as word embeddings are exposed to controlled noise since jamming cannot be mitigated perfectly. This exposure further helps improve the model robustness by teaching it to learn effectively even in the presence of interference [24].
@@ II System Model and Adversarial Analysis
II-A Wireless R-SFLLM System Model
We consider an SFL setup in which a set $\mathcal{Q}$ of $Q$ legitimate clients cooperatively train transformer-based LLMs, which consist of embedding, attention and head layers. A natural choice in SFL with LLMs is to partition the model according to these blocks, assigning the embedding layer to the client and the compute-intensive attention and head layers to the server as outlined in Figure 1. This particular partitioning alleviates the computational load at the client while ensuring that word embeddings are processed close to the raw data, thereby enhancing privacy. Further partitioning the embedding block and transmitting intermediate layers instead increases the risk of sensitive information being exposed to adversarial attacks, such as model inversion [25]. During training, each user $q\in\mathcal{Q}$ first computes the word embeddings $\boldsymbol{e}_{q}\in\mathbb{R}^{E}$ for its private data points and then maps the embeddings onto uncorrelated zero-mean, unit variance Gaussian symbols, which are then beamformed, power-scaled, and transmitted over the wireless channel to a dedicated server slice for further processing.
Figure 1: SFL model split with LLM word embeddings being processed at the client and with attention and head layers being processed at the server.
To this end, we consider a MIMO-OFDM multiple access channel (MAC) in the uplink. Each user $q$ is equipped with $N_{T_{q}}$ antennas and transmits the signal $\boldsymbol{x}_{qnk}$, which is a composite of the binary user scheduling $\alpha_{qnk}$, transmit power $p_{qnk}$, beamforming vector $\boldsymbol{w}_{qnk}\in\mathbb{C}^{N_{T_{q}}}$, and word embedding data symbols $s_{qnk}$. The transmissions occur over the resource set $\mathcal{R}_{q}=\mathcal{N}_{q}\times\mathcal{K}_{q}$ with allocated subcarriers $n\in\mathcal{N}_{q}$ and OFDM symbols $k\in\mathcal{K}_{q}$, with a total of $N$ subcarriers and $K$ symbols available. Each legitimate transmit signal propagates through the channel $\boldsymbol{H}_{qnk}\in\mathbb{C}^{N_{R}\times N_{T_{q}}}$ to the server, equipped with $N_{R}$ receive antennas. In this setup, an adversarial jammer aims to impair the SFL training by jamming the word embeddings in the uplink. The legitimate signal is thus corrupted by additive white Gaussian noise (AWGN) $\boldsymbol{\eta}_{nk}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I})\in\mathbb{C}^{N_{R}}$ and by the adversarial jamming signal $\boldsymbol{u}_{nk}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{C}_{\boldsymbol{u}_{nk}})\in\mathbb{C}^{N_{J}}$, which propagates through the separate jamming channel $\boldsymbol{G}_{nk}\in\mathbb{C}^{N_{R}\times N_{J}}$. We define $N_{J}$ as the number of jamming antennas and corresponding jamming covariance matrix as $\boldsymbol{C}_{\boldsymbol{u}_{nk}}\in\mathbb{C}^{N_{J}\times N_{J}}$. The receiver performs equalization using the linear filters $\boldsymbol{v}_{qnk}^{H}\in\mathbb{C}^{1\times N_{R}}$ to estimate the transmitted symbols $\hat{s}_{qnk}$. In summary, we have:
@@ II-A Wireless R-SFLLM System Model
Here, $L_{H_{q}}$ and $L_{G}$ are the number of resolvable paths for each channel, $b_{\boldsymbol{\cdot},l}$ is the path gain for each resolvable path $l$, $\boldsymbol{a}_{N_{X}}(\boldsymbol{\theta})$ is the steering vector at each terminal with $N_{X}$ antennas, $\boldsymbol{\theta}_{\boldsymbol{\cdot},l}$ is direction-of-arrival, $\boldsymbol{\psi}_{\boldsymbol{\cdot},l}$ is direction-of-departure, and $w_{nk}(\nu,\tau)=k\nu T_{s}-n\tau\Delta f$ is the phase shift caused by the Doppler shift $\nu$ and propagation delay $\tau$, with $T_{s}$ and $\Delta f$ being the symbol period and subcarrier spacing. Furthermore, the power $P_{q}$ for each user is limited across all resource elements and the jamming signal equally adheres to a jamming power constraint $P_{J}$, i.e.
@@
We further assume that the legitimate parties have precise channel state information (CSI), encompassing the wireless link parameters defined by the set $\zeta_{q}=\{\alpha_{qnk},p_{qnk},\boldsymbol{w}_{qnk},\boldsymbol{v}_{qnk},\boldsymbol{H}_{qnk},\sigma^{2}\}$. Additionally, the SFL participants are provided with the DoAs of the jamming signal, i.e. $\boldsymbol{\theta}_{G}=\{\theta_{G,l}\}_{l=1}^{L_{G}}$. This may be enabled by advanced wireless sensing technologies in future 6G networks, such as integrated sensing and communication (ISAC) and reflective intelligent surfaces (RIS) [26, 27, 28], to name a few. Further, we assume no restrictions on the particular jamming strategy, hence the jammer is assumed to be in the so-called jammer-dominant regime [20] with more transmit power and antennas than any legitimate party, i.e. $P_{J}\gg P_{q}$ and $N_{J}>N_{T_{q}},N_{R}$. In addition, the adversary may possess full system knowledge, including $\zeta_{q}$. This represents a worst-case jammer assumption.
The adversarial jamming introduces corruption not only at the symbol level but also at the decoded message, such that jamming can be modeled as the post-decoding error as follows:
@@
Upon receiving the jammed signal $\boldsymbol{y}_{nk}$, the server slice continues processing the LLM attention and head layers using the corrupted word embeddings $\hat{\boldsymbol{e}}_{q}\neq\boldsymbol{e}_{q}$. At the end of the forward propagation pass, the server slice computes the training loss metric $L:\mathbb{R}^{E}\rightarrow\mathbb{R}$, which yields the corrupted loss $L(\hat{\boldsymbol{e}}_{q})$ and its gradient $\nabla L(\hat{\boldsymbol{e}}_{q})$, using which the backpropagation process is initiated. This procedure is repeated for each transmission of the word embeddings across all global training rounds. We also assume that the jammer is not active in the downlink as the perturbation of gradients has been studied in various federated setups, for which corresponding defense mechanisms exist [29]. Similarly, the client- and server-side model aggregation after each global round are assumed to be unaffected by the adversary as corresponding secure aggregation strategies exist as well [30]. Note that FedAvg [9] is used in this work. Thus, the consideration of only the uplink transmission suffices to study the jamming impact on LLM word embeddings in this setup. Anti-jamming can then be directly applied if necessary conditions are fulfilled, including the assumption that maximizing the signal-to-interference-plus-noise-ratio (SINR) implies maximizing the ML performance. This assumption is verified next. Figure 2 shows the R-SFLLM system architecture, where the SFL protocol is augmented by sensing-assisted jamming DoA information, a necessary component for our anti-jamming framework in Section III.
II-B Adversarial Jamming Impact on LLM Training in SFL
Previous works in FL typically assume the loss function to be convex, twice differentiable, and Lipschitz smooth. While these assumptions may hold true for simpler neural networks as in [12] and [31], more involved architectures such as transfomers generally do not exhibit these properties [32]. The assumption that $L$ is Lipschitz smooth is particularly far-reaching as this implies bounded gradients during backpropagation. In [33], it is shown that the standard Lipschitz assumption introduces a large variability along the optimization trajectory. Thus, a relaxed $(\boldsymbol{L_{0}},\boldsymbol{L_{1}})$-smoothness needs to be assumed, which generalizes to more complex models, such as LLMs. Based on this generalization, we derive upper bounds on the loss divergence, caused by jammed word embeddings, and show how these relate to the communication MSE.

### v3

@@ Abstract
Split federated learning (SFL) is a compute-efficient paradigm in distributed machine learning (ML), where components of large ML models are outsourced to remote servers. A significant challenge in SFL, particularly when deployed over wireless channels, is the susceptibility of transmitted model parameters to adversarial jamming that could jeopardize the learning process. This is particularly pronounced for embedding parameters in large language models (LLMs) and vision language models (VLMs), which are learned feature vectors essential for domain understanding. In this paper, rigorous insights are provided into the influence of jamming embeddings in SFL by deriving an expression for the ML training loss divergence and showing that it is upper-bounded by the mean squared error (MSE). Based on this analysis, a physical layer framework is developed for resilient SFL with LLMs (R-SFLLM [^1] Code available at: https://github.com/aladinD/R_SFLLM) over wireless networks. R-SFLLM leverages wireless sensing data to gather information on the jamming directions-of-arrival (DoAs) for the purpose of devising a novel, sensing-assisted anti-jamming strategy while jointly optimizing beamforming, user scheduling, and resource allocation. Extensive experiments using both LLMs and VLMs demonstrate R-SFLLM’s effectiveness, achieving close-to-baseline performance across various natural language processing (NLP) and computer vision (CV) tasks, datasets, and modalities. The proposed methodology further introduces an adversarial training component, where controlled noise exposure significantly enhances the model’s resilience to perturbed parameters during training. The results show that more noise-sensitive models, such as RoBERTa, benefit from this feature, especially when resource allocation is unfair. It is also shown that worst-case jamming in particular translates into worst-case model outcomes, thereby necessitating the need for jamming-resilient SFL protocols.
@@ I Introduction and Motivation
Future 6G networks are anticipated to introduce a substantial leap toward highly integrated and intelligent connectivity at the edge, enabled by artificial intelligence (AI) [1] and machine learning (ML) [2]. However, in this envisioned hyper-connected and AI-assisted network, important questions and stringent requirements on resilience and trustworthiness arise, both from a user and network perspective [3]. This imposes significant design challenges for emerging technologies, such as distributed and collaborative ML (DCML) [4], which may be targeted by adversarial attacks over the wireless medium. For instance, the distributed training of large language models (LLMs) faces unique challenges in ensuring data integrity and model resilience due to highly sensitive embedding parameters. Much research has focused on addressing these challenges via adversarial AI training methods [5] and model-based solutions [6], and in [7], the authors closely investigated such attack and defense strategies, particularly in 6G networks, exposing critical vulnerabilities across all network layers. However, a number of important research questions remain underexplored, such as how these attacks can be orchestrated in practice, and how current and next-generation wireless network architectures, systems, and technologies can introduce proactive defense mechanisms, ideally by-design.
I-A Adversarial Poisoning in Wireless Federated LLM Training.
@@
Motivated by the increasing importance of end-user data privacy and associated privacy protection laws [8], DCML has been gradually shifting toward mechanisms such as federated learning (FL) [9] and split FL (SFL) [10]. These have proven to be privacy-preserving as only the respective model parameters, some parts of them, or model gradients need to be exchanged. SFL in particular has emerged as a compute-efficient federated protocol, suitable for distributed training of large ML model architectures. Unlike traditional FL, in which the entire model is trained on each client, SFL splits the model, allowing for more compute-intensive parts to be outsourced to a remote server. This approach is particularly advantageous for LLM architectures that cannot be entirely processed at resource-constrained edge devices due to computational and memory limitations [11]. However, the practical realization of (S)FL over wireless networks faces challenges from the inherently unreliable wireless medium, limited bandwidth, and suboptimal resource allocation [12]. In addition, malicious actors such as jammers could target privacy and security aspects of (S)FL systems by intentionally poisoning data and models through adversarial noise. The particular importance of studying adversarial jamming attacks in SFL with LLMs is motivated by recent results from natural language processing (NLP) research [13, 14]. The authors in [13] study the susceptibility of LLMs to word embedding poisoning caused by noisy perturbations and show that by altering even a single embedding vector, an adversary can subtly manipulate a model to react abnormally to specific trigger words. Moreover, the severity of such embedding poisoning attacks for federated networks was studied in [14], showing that even a small number of compromised clients is sufficient to effectively deteriorate the global model. In federated systems over wireless networks, adversarial jamming emerges as a realistic threat for orchestrating such poisoning attacks by corrupting sensitive word or other multi-modal embeddings during transmission. This covert adversarial perspective is in contrast to prior works on jamming in FL [15, 16], which do not consider the implications of poisoning the model’s reasoning capabilities. This is particularly critical in SFL, where embeddings might be directly transmitted as intermediate split parameters. However, the detrimental impact of such attacks in SFL remains underexplored.
I-B Proactive and Resilient-by-Design Anti-Jamming in SFL.
@@
To preemptively safeguard LLM parameter transmissions in SFL against adversarial jamming attacks, proactive and in particular resilient-by-design approaches are needed [17]. This requires a simultaneous co-design of AI, resilience, beamforming, user scheduling, and resource allocation, thereby integrating resilience from a bottom-up approach, by design. In addition, such proactive defense mechanisms need to be agnostic of the respective jamming capabilities, including physical and spatial features. This is not the case for some more recent works, which tend to impose strong assumptions on the adversary’s knowledge and setup. For instance, [18] and [19] only consider single- or few-antenna jammers with common secrets being exchanged between legitimate parties. This essentially excludes so-called worst-case jammers with extensive system knowledge and capabilities [20, 21]. To develop universally applicable defense strategies for a wide range of jamming scenarios in SFL, system performance needs to be guaranteed for the worst-case. Thus, we need to generalize toward intelligent and reconfigurable worst-case jammers. In our prior work in [22, 23], we studied how sensing-assisted network information can be harnessed to enhance existing mitigation schemes without the need for otherwise precise jamming statistics. Therein, we have shown that information on the jamming signal directions-of-arrival (DoAs) can be used to devise MIMO-OFDM anti-jamming strategies with exceptional performance. However, we did not discuss whether such sensing-assisted defense strategies can be straightforwardly applied to enhance the resilience in SFL over wireless networks. In particular, the impact of worst-case jamming needs to be quantified in order to study its influence on LLM model poisoning as compared to conventional jammers. In the subsequent sections, we provide thorough insights on these aspects, including an analysis on the minimum system rate that guarantees a reliable and resilient SFL training.
I-C Contributions.
@@
The main contribution of this paper is an analysis and framework for resilient SFL over wireless networks that will help close the gap between adversarial jamming attacks in SFL and LLM model poisoning. Here, resilience transcends conventional robustness by proactively integrating physical-layer defenses, ensuring the global training process remains effective even under worst-case jamming, rather than merely withstanding moderate perturbations. We provide insights into how jamming of LLM embeddings affects the global model training and how the latter can be effectively safeguarded by MIMO signal processing at the availability of sensing-assisted DoA information. In summary, our key contributions include:
  1. 1.
@@
    Analytical Bound for LLM Loss Divergence: We derive a novel expression for the training loss divergence in case of corrupted embeddings under a relaxed $(\boldsymbol{L}_{0},\boldsymbol{L}_{1})$-smoothness assumption, and show that its upper bound depends on the communication mean squared error (MSE), thereby motivating a wireless approach to resilience in SFL.
  2. 2.
    Minimum System Rate Guarantees for SFL: We provide a novel analysis on the minimum system rate that ensures reliable SFL over wireless networks. By relating outage rates to jamming power and user scheduling, we characterize key scalability conditions and illustrate how larger-scale SFL remains feasible under adversarial interference.
  3. 3.
@@
    Sensing-Assisted Anti-Jamming Framework: We develop R-SFLLM, a novel, sensing-assisted anti-jamming framework for resilient SFL with LLMs, which leverages the jammer’s DoAs to devise an anti-jamming strategy formulated as a joint optimization problem for beamforming, user scheduling, and resource allocation while maximizing the sum rate of the SFL participants. In this problem, any explicit knowledge about the jamming statistics is replaced by a surrogate expression that depends only on the jamming DoAs. We provide an efficient solution to the problem using an iterative water-filling approach [24], thereby going beyond mere algorithmic robustness and enabling resilience-by-design through the physical layer.
  4. 4.
    Worst-Case Jamming: To validate R-SFLLM under severe threat conditions, we develop the worst-case jamming strategy [22], which minimizes the total system sum rate.
  5. 5.
@@
    Extensive Experiments: We provide results for NLP-based BERT [25] and RoBERTa [26] LLMs, as well as computer vision (CV)-based CLIP [27] vision language models (VLMs), covering various tasks across 13 datasets and two modalities. We demonstrate near-optimal performance when anti-jamming is enabled and significantly worse outcomes for unprotected scenarios. Further, we show that R-SFLLM introduces an adversarial training component as embeddings are exposed to controlled noise since jamming cannot be mitigated perfectly, thus improving the model’s resilience by teaching it to learn effectively in the presence of interference [28]. We also provide ablation studies on scalability and jamming capabilities.
@@ II System Model and Adversarial Analysis
II-A Wireless R-SFLLM System Model.
We consider an SFL setup in which a set $\mathcal{Q}$ of $Q$ legitimate clients cooperatively train transformer-based LLMs (analogously for VLMs), which consist of embedding, attention, and head layers. A natural choice in SFL is to partition the model according to these blocks, assigning the embedding layer to the client and the compute-intensive attention and head layers to the server, as shown in Fig. 1. This particular partitioning alleviates the computational load at the client while ensuring that embeddings are processed close to the raw data, thereby enhancing privacy. Further partitioning the embedding block and transmitting intermediate layers instead increases the risk of sensitive information being exposed to adversarial attacks, such as model inversion [29]. During training, each user $q\in\mathcal{Q}$ first computes the embeddings $\boldsymbol{e}_{q}\in\mathbb{R}^{E}$ for its private data points and then maps them onto uncorrelated zero-mean, unit variance Gaussian symbols, which are then beamformed, power-scaled, and transmitted over the wireless channel to a dedicated server slice for further processing. To this end, we consider a MIMO-OFDM multiple access channel (MAC) in the uplink. Each user $q$ is equipped with $N_{T_{q}}$ antennas and transmits the signal $\boldsymbol{x}_{qnk}$, which is a composite of the binary user scheduling $\alpha_{qnk}$, transmit power $p_{qnk}$, beamforming vector $\boldsymbol{w}_{qnk}\in\mathbb{C}^{N_{T_{q}}}$, and embedding data symbols $s_{qnk}$. The transmissions occur over the resource set $\mathcal{R}_{q}=\mathcal{N}_{q}\times\mathcal{K}_{q}$ with allocated subcarriers $n\in\mathcal{N}_{q}$ and OFDM symbols $k\in\mathcal{K}_{q}$, with a total of $N$ subcarriers and $K$ symbols available. Each legitimate transmit signal propagates through the channel $\boldsymbol{H}_{qnk}\in\mathbb{C}^{N_{R}\times N_{T_{q}}}$ to the server with $N_{R}$ receive antennas. An adversarial jammer aims to impair the SFL training by jamming the embeddings in the uplink. The legitimate signal is thus corrupted by additive white Gaussian noise (AWGN) $\boldsymbol{\eta}_{nk}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I})\in\mathbb{C}^{N_{R}}$ and by the adversarial jamming signal $\boldsymbol{u}_{nk}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{C}_{\boldsymbol{u}_{nk}})\in\mathbb{C}^{N_{J}}$, which propagates through the separate jamming channel $\boldsymbol{G}_{nk}\in\mathbb{C}^{N_{R}\times N_{J}}$. We define $N_{J}$ as the number of jamming antennas and corresponding jamming covariance matrix as $\boldsymbol{C}_{\boldsymbol{u}_{nk}}\in\mathbb{C}^{N_{J}\times N_{J}}$. The receiver performs equalization using the linear filters $\boldsymbol{v}_{qnk}^{H}\in\mathbb{C}^{1\times N_{R}}$ to estimate the transmitted symbols $\hat{s}_{qnk}$. In summary, we have:
@@
Figure 1: SFL model split with LLM embeddings being processed at the client and with attention and head layers being processed at the server.
Here, $L_{H_{q}}$ and $L_{G}$ are the number of resolvable paths $l$ for each channel, $b_{\boldsymbol{\cdot},l}$ is the path gain, $\boldsymbol{a}_{N_{X}}(\boldsymbol{\theta})$ is the steering vector at each terminal with $N_{X}$ antennas, $\boldsymbol{\theta}_{\boldsymbol{\cdot},l}$ is direction-of-arrival, $\boldsymbol{\psi}_{\boldsymbol{\cdot},l}$ is direction-of-departure, and $w_{nk}(\nu,\tau)=k\nu T_{s}-n\tau\Delta f$ is the phase shift caused by the Doppler shift $\nu$ and propagation delay $\tau$, with $T_{s}$ and $\Delta f$ being the symbol period and subcarrier spacing. Furthermore, the power $P_{q}$ for each user is limited across all resource elements, and the jamming signal equally adheres to a jamming power constraint $P_{J}$, i.e.
@@
We further assume that the legitimate parties have precise channel state information (CSI) (e.g., via pilot‑echoes), encompassing the wireless link parameters defined by the set $\zeta_{q}=\{\alpha_{qnk},p_{qnk},\boldsymbol{w}_{qnk},\boldsymbol{v}_{qnk},\boldsymbol{H}_{qnk},\sigma^{2}\}$. Additionally, the SFL participants are provided with the DoAs of the jamming signal, i.e. $\boldsymbol{\theta}_{G}=\{\theta_{G,l}\}_{l=1}^{L_{G}}$. This may be enabled by advanced wireless sensing technologies in future 6G networks, such as integrated sensing and communication (ISAC) and reflective intelligent surfaces (RIS) [30, 31, 32, 33, 34], to name a few, thus making it a realistic assumption in practice. Further, we assume no restrictions on the particular jamming strategy, hence the jammer is assumed to be in the so-called jammer-dominant regime [20] with more transmit power and antennas than any legitimate party, i.e. $P_{J}\gg P_{q}$ and $N_{J}>N_{T_{q}},N_{R}$. In addition, the adversary may possess full system knowledge, including $\zeta_{q}$. This represents a worst-case jammer assumption.
Adversarial jamming introduces corruption not only at the symbol level but also at the decoded message, such that it can be modeled as the post-decoding error as follows:
@@
Upon receiving the jammed signal $\boldsymbol{y}_{nk}$, the server continues processing the LLM attention and head layers using the corrupted embeddings $\hat{\boldsymbol{e}}_{q}\neq\boldsymbol{e}_{q}$. At the end of the forward propagation pass, the server computes the training loss $L:\mathbb{R}^{E}\rightarrow\mathbb{R}$, which yields the corrupted loss $L(\hat{\boldsymbol{e}}_{q})$ and its gradient $\nabla L(\hat{\boldsymbol{e}}_{q})$, using which the backpropagation process is initiated. This procedure is repeated for each transmission of the embeddings across all global training rounds. We also assume that the jammer is not active in the downlink as the perturbation of gradients has been studied in various federated setups, for which corresponding defense mechanisms exist [35]. Similarly, the client- and server-side model aggregation after each global round are assumed to be unaffected by the adversary as corresponding secure aggregation strategies exist as well [36]. Note that FedAvg [9] is used in this work. Thus, the consideration of only the uplink transmission suffices to study the impact of jamming LLM embeddings in this setup. Anti-jamming can then be directly applied if necessary conditions are fulfilled, including the assumption that maximizing the signal-to-interference-plus-noise-ratio (SINR) implies maximizing the ML performance. This assumption is verified next. Fig. 2 shows the R-SFLLM system architecture, where the SFL protocol is augmented by sensing-assisted jamming DoA information, a necessary component for our anti-jamming framework presented in Section III.
II-B Adversarial Jamming Impact on LLM Training in SFL.
Previous works in (S)FL typically assume the loss function to be convex, twice differentiable, and Lipschitz smooth. While these assumptions may hold true for simpler neural networks as in [12] and [37], more involved architectures such as transfomers generally do not exhibit these properties [38]. The assumption that $L$ is Lipschitz smooth is particularly far-reaching as this implies bounded gradients during backpropagation. In [39], it is shown that the standard Lipschitz assumption introduces a large variability along the optimization trajectory. Thus, a relaxed $(\boldsymbol{L_{0}},\boldsymbol{L_{1}})$-smoothness needs to be assumed, which generalizes to more complex models, such as LLMs/VLMs. Based on this generalization, we derive upper bounds on the loss divergence, caused by jammed embeddings, and show how these relate to the communication MSE.